# Cosmic superstrings, metastable strings and ultralight primordial black holes: from NANOGrav to LIGO and beyond

### v1

@@ IV.1 Complementary GW signatures: density fluctuations and graviton emission
Figure 6: Top-left: Combined GW spectrum (Metastable strings+density fluctuation) for the benchmark parameters belonging to Category C, see Table 2. The spectrum has been produced for $\beta=5\times 10^{-9}$ and summing over $10^{3}$ modes of cosmic string loop oscillations. Top-right: GWs only from metastable cosmic strings for the same benchmark parameters but for different values of $\beta$, showing the impact of the finite duration of PBH domination. Bottom: GW strain $h_{c}$ from density fluctuation (dashed), from Metastable strings (solid), and graviton radiation from PBH (dotted) with the projected sensitivities (including the black curly curve) of various high-frequency GW detectors. All the benchmark parameters are the same as in the top-left plot. The region above the red dashed line is excluded from $\Delta N_{\rm eff}$ bound by Planck.
@@
Figure 7: Top-left: Combined GW spectrum (Metastable strings+density fluctuation) for the benchmark parameters belonging to Category C, see Table 2. The spectrum has been produced for $\beta=3\times 10^{-8}$ and summing over $10^{3}$ modes of cosmic string loop oscillations. Top-right: GWs only from metastable cosmic strings for the same benchmark parameters but for different values of $\beta$, showing the impact of the finite duration of PBH domination. Bottom: GW strain $h_{c}$ from density fluctuation (dashed), from Metastable strings (solid), and graviton radiation from PBH (dotted) with the projected sensitivities (including the black curly curve) of various high-frequency GW detectors. All the benchmark parameters are the same as in the top-left plot. The region above the red dashed line is excluded from $\Delta N_{\rm eff}$ bound by Planck.
@@
The ultralight PBHs are many-faced objects concerning the stochastic GWs. For instance, the initial curvature perturbations leading to the formation of PBHs give rise to GWs (see, e.g., [128, 129, 130]), ultralight PBHs could possibly merge, releasing GWs [131], PBHs emit gravitons, which constitute GWs at ultra-high frequencies [6, 7, 8, 9], and the inhomogeneous distribution of PBHs, which results in density fluctuations, triggers the production of GWs [10, 11, 12]. Here we do not discuss any specific production mechanisms of PBHs, nor the possibility of merger formation. We rather focus on the last two options, being slightly biased to the fact that graviton emission is inevitable and independent of the production mechanism, the GWs from PBH-density fluctuation seem to provide a realistic stochastic background with spectral shape dependent on their spatial distribution which has been assumed to be like poison like in Refs.[10, 11]. Even though the PBH gas, on average, behaves like pressure-less dust, the inhomogeneous spatial distribution results in density fluctuations, which are isocurvature in nature. When PBHs dominate the universe’s energy density, the isocurvature component gets converted to curvature perturbation, which subsequently induces secondary GWs. Due to the significant density fluctuations at smaller scales (equivalent to the average separation of PBHs at $T_{\rm Bf}$), substantial GWs are generated. These GWs are further amplified due to the nearly instantaneous evaporation of PBHs. The amplitude of such induced GWs in the present day is given by
@@
The ultra-violet cutoff frequency $f_{\rm peak}$ corresponds to the comoving scale representing the mean separation of PBHs at the formation time.
@@
In Fig.6 and Fig.7 (top-left) we show the total GW spectrum: density fluctuation spectrum combined with the one from metastable strings and superstrings for the benchmark parameters belonging to Category C in each string class. Instead of large $\beta$, here we chose a benchmark value such that the duration of PBH domination is relatively short and the density fluctuation GW spectrum is visible (the sharp peak at higher frequencies). Note that because of the strong dependence of the peak GW amplitude on $\beta$ (see Eq.(IV.2)) in this scenario, the duration of PBH domination can not be arbitrarily large, otherwise, the GWs would saturate the BBN bound on $\Delta N_{\rm eff}$: $\int{\rm d}f~{}f^{-1}\Omega_{\rm GW}(f)h^{2}\lesssim 5.6\times 10^{-6}\Delta N_{\rm eff}$, with $\Delta N_{\rm eff}\lesssim 0.2$. Consequently, the falling of the GW spectrum from cosmic strings would stop at higher frequencies, potentially showing up a subsequent GW plateau originating from the loops in the radiation epoch before the PBH domination. This has been shown in the top-right panel of Fig.6 for metastable strings for three benchmark values of $\beta$. Nonetheless, the appearance of the second plateau is not always the case if the particle production cut-off is strong enough. We show it for superstrings on the top-right panel of Fig.7. Compared to the metastable string case, here owing to the strong particle production cut-off, the GW spectra closely follow the large $\beta$ case. As discussed in Sec.II, for the GWs with amplitude at the level of PTAs, the particle production cut-off appears mostly at frequencies higher than the LIGO frequency band. Therefore, for shorter-duration PBH-domination scenarios, the regions in Fig.3 and Fig.4 extracted for $\Omega_{\rm GW}\propto f^{-1/3}$ falling, requires modification now to account for a flattening of the spectrum. Although we do not present such a scan, to observe GW peak from density fluctuation on top of the strong amplitude cosmic string radiated GWs required by PTAs, reasonably large $\beta$ (respecting, of course, the BBN bound) is required. For this reason, for most of the parameter space presented in Fig.3 and Fig.4, the duration of PBH domination would be such that from $f_{\rm brk}$ until the $f_{\rm LIGO}$, the spectrum would fall as $\Omega_{\rm GW}\propto f^{-1/3}$, keeping the results of the previous section intact. We expect modification to the parameter space for heavier PBH masses around $10^{8}$ g leading to smaller values of $f_{\rm brk}$, and thereby potential flatting of the spectrum at the LIGO frequency band due to the shorter duration of PBH domination.
Besides these low-frequency peculiarities, a PBH+superstrings (metastable strings) system provides a characteristic high-frequency GW spectrum constituting gravitons as shown in the bottom panels of Fig.6 and 7, where the GW strain $h_{c}$ is related to $\Omega_{\rm GW}$ as
@@
Because Eq.(IV.5) characterizes the GW background, for the allowed mass ranges of PBH presented in Sec.III, a considerable GW background constituting gravitons can be obtained within a constrained frequency range spanning approximately, only a decade. Although the proposed high-frequency GW detectors [132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142] are not presently competitive with the $\Delta N_{\rm eff}$ bound from Planck [143] (according to Ref.[144], the electromagnetic resonant cavities can potentially reach below the dark radiation projections around $f\sim 10^{8}$ Hz; the black oscillating curve in Fig.6 and Fig.7), optimistically, in the future graviton background may be detected. In any case, these marked spectral features in the total GW background are described with three characteristic frequencies $f_{\rm brk}$, $f_{\rm peak,I}$, and $f_{\rm peak,g}$ make the PBH + string scenario more appealing than an ordinary EMD+string case.
@@ IV.2 Implication on a realistic BSM model of leptogenesis and super-heavy dark matter
where we have assumed leptogenesis is dominated by the decay of $N_{1}$ so that $M_{1}\sim T_{\rm lepto}$. Eq.(IV.8) can be shown to be accurate, matching the numerical results obtained from Eq.(III.1),(III.2) and the modified leptogenesis equations in the presence of PBHs [49]. The PBHs can also populate the decaying RHNs to produce lepton asymmetry, but if Eq.(IV.7) is to be satisfied, such PBHs evaporate after Spaleron freeze-out temperature, and the lepton asymmetry does not get processed to BAU. Therefore we discard this option.
@@
$\mu=\pi M_{DM}^{2}~{}f_{\rm DM}^{-2}h\left(\lambda,g^{\prime}\right),$ (IV.9)
@@
With this relation, we now redo the PTA fit for metastable strings performed in the previous section. The result is shown in Fig.9. Note that because of this extra condition in Eq.(IV.10), this $U(1)_{B-L}^{\rm PBH}$ model fits the PTA data for an extremely constrained range of PBH mass and produces GW signals belonging only to Category A. Therefore any non-observation of SGWB in the LVK-D would rule out the model/the best parameter choice of the model. Besides, because the PBH mass window is narrow, the other GW signals, e.g., from density fluctuation plus in the form of gravitons are peaked in a stringent range in frequencies. In an upcoming publication, we shall provide a detailed discussion of the model for more general parameter space with all possible GW signatures.
@@ V Conclusion
An abundant presence of ultralight PBHs leading to their domination on energy densities in the early universe could produce plenty of observable signatures specifically in the form of gravitational radiation. In addition to producing direct GW background, e.g., in the form of gravitons, a PBH domination gets imprinted on the spectral shape of GWs originating from other independent sources. Such a spectral shape reconstruction with the GW detectors and studying complementary GW signals originating directly from the PBHs could be an interesting way to probe the properties of ultralight PBHs. Here we study the impact of PBH domination on the GWs emitted from cosmic superstrings and metastable strings which are speculated to be among the sources of nHz GWs observed by the PTA. Such an interpretation might become perilous, as a GW background originating from cosmic strings spans a wide range in frequencies thereby contradicting the non-observation of stochastic GW background by LIGO dealing with comparable signal strength. In fact, the LIGO constraint rules out a significant string-parameter space favoured by the recent PTA data. A PBH-dominated early universe, however, affects the propagation of such GWs to reduce the amplitude at the high frequencies thereby evading the LIGO constraints. Assuming monochromatic and non-rotating PBHs we work out the mass range of such ultralight PBHs which makes the string-radiated GWs consistent with LIGO bound and the PTA data. We show that PBHs with mass in the range $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$ would produce such viable signals that can be tested in the mid-band detectors such as LISA, high-frequency detectors such as LVK-D plus ET with a high signal-to-noise ratio. We discuss why a PBH-domination is distinguishable from any other early matter-domination scenarios that reduce the amplitude of string-radiated GWs at higher frequencies. In this context, we discuss a BSM model based on $U(1)_{B-L}$-seesaw providing neutrino mass, leptogenesis, super-heavy dark matter, and metastable cosmic strings. We show how a PBH-cosmic string scenario could be extremely predictive in probing unexplored parameter space of such BSM scenarios.
@@ Appendix A Superstrings and Metastable strings fit to the NANOGrav data
Figure 10: Triangular plot on the left (right) from the MCMC runs for superstrings (metastable strings) with the NANOGrav 15-year dataset using PTArcade[153].

### v2

@@ IV.1 Complementary GW signatures: density fluctuations and graviton emission
Figure 7: Top-left: Combined GW spectrum (Metastable strings + density fluctuations) for the benchmark parameters belonging to Category C, see Table 2. The spectrum has been produced for $\beta=10^{-8}$. Top-right: GWs only from metastable cosmic strings for the same benchmark parameters but for different values of $\beta$, showing the impact of the finite duration of PBH domination. Bottom: GW strain $h_{c}$ from density fluctuations (dashed), from Metastable strings (solid), and graviton radiation from PBH (dotted) with the projected sensitivities (including the black curly curve) of various high-frequency GW detectors. All the benchmark parameters are the same as in the top-left plot. The region above the red dashed line is excluded from $\Delta N_{\rm eff}$ bound by Planck.
@@
Figure 8: Top-left: Combined GW spectrum (Superstrings + density fluctuations) for the benchmark parameters belonging to Category C, see Table 2. The spectrum has been produced for $\beta=8\times 10^{-9}$. Top-right: GWs only from superstrings for the same benchmark parameters but for different values of $\beta$, showing the impact of the finite duration of PBH domination. Bottom: GW strain $h_{c}$ from density fluctuations (dashed), from superstrings (solid), and graviton radiation from PBH (dotted) with the projected sensitivities (including the black curly curve) of various high-frequency GW detectors. All the benchmark parameters are the same as in the top-left plot. The region above the red dashed line is excluded from $\Delta N_{\rm eff}$ bound by Planck.
@@
The ultralight PBHs are many-faced objects concerning the stochastic GWs. For instance, the initial curvature perturbations leading to the formation of PBHs give rise to GWs (see, e.g., [138, 139, 140, 141]), ultralight PBHs could possibly merge, releasing GWs [142], PBHs emit gravitons, which constitute GWs at ultra-high frequencies [6, 7, 8, 9], and the inhomogeneous distribution of PBHs, which results in density fluctuations, triggers the production of GWs [10, 11, 12]. Here we do not discuss any specific production mechanisms of PBHs, nor the possibility of merger formation. We rather focus on the last two options, being slightly biased to the fact that graviton emission is inevitable and independent of the production mechanism, the GWs from PBH-density fluctuations seem to provide a realistic stochastic background with spectral shape dependent on their spatial distribution which has been assumed to be like Poisson like in Refs.[10, 11]. Even though the PBH gas, on average, behaves like pressure-less dust, the inhomogeneous spatial distribution results in density fluctuations, which are isocurvature in nature. When PBHs dominate the universe’s energy density, the isocurvature component gets converted to curvature perturbation, which subsequently induces secondary GWs. Due to the significant density fluctuations at smaller scales (equivalent to the average separation of PBHs at $T_{\rm Bf}$), substantial GWs are generated. These GWs are further amplified due to the nearly instantaneous evaporation of PBHs. The amplitude of such induced GWs in the present day is given by [14]
@@
The ultra-violet cutoff frequency $f_{\rm peak,I}$ corresponds to the comoving scale representing the mean separation of PBHs at the formation time.
@@
In Fig.7 and Fig.8 (top-left) we show the total GW spectrum: density fluctuations spectrum combined with the one from metastable strings and superstrings for the benchmark parameters belonging to Category C in each string class. Instead of large $\beta$, here we chose a benchmark value such that the duration of PBH domination is relatively short and the density fluctuations GW spectrum is visible (the sharp peak at higher frequencies). Note that because of the strong dependence of the peak GW amplitude on $\beta$ (see Eq.(IV.2)) in this scenario, the duration of PBH domination can not be arbitrarily large, otherwise, the GWs would saturate the BBN bound on $\Delta N_{\rm eff}$: $\int{\rm d}f~{}f^{-1}\Omega_{\rm GW}(f)h^{2}\lesssim 5.6\times 10^{-6}\Delta N_{\rm eff}$, with $\Delta N_{\rm eff}\lesssim 0.2$. Consequently, the falling of the GW spectrum from cosmic strings would stop at higher frequencies, potentially showing up a subsequent GW plateau originating from the loops in the radiation epoch before the PBH domination. This has been shown in the top-right panel of Fig.7 for metastable strings for three benchmark values of $\beta$. Nonetheless, the appearance of the second plateau is not always the case if the particle production cut-off is strong enough. We show it for superstrings on the top-right panel of Fig.8. Compared to the metastable string case, here owing to the strong particle production cut-off, the GW spectra closely follow the large $\beta$ case. As discussed in Sec.II, for the GWs with amplitude at the level of PTAs, the particle production cut-off appears mostly at frequencies higher than the LIGO frequency band. Therefore, for extremely shorter-duration PBH-domination scenarios, the regions in Fig.4 and Fig.5 extracted for $\Omega_{\rm GW}\propto f^{-1/3}$ falling, requires modification now to account for a flattening of the spectrum. Although we do not present such a scan, to observe GW peak from density fluctuations on top of the strong amplitude cosmic string radiated GWs required by PTAs, reasonably large $\beta$ (respecting, of course, the BBN bound) is required. For this reason, for most of the parameter space presented in Fig.4 and Fig.5, the duration of PBH domination would be such that from $f_{\rm brk}$ until the $f_{\rm LIGO}$, the spectrum would fall as $\Omega_{\rm GW}\propto f^{-1/3}$, keeping the results of the previous section intact. We expect modification to the parameter space for heavier PBH masses around $10^{8}$ g leading to smaller values of $f_{\rm brk}$, and thereby potential flatting of the spectrum at the LIGO frequency band due to the shorter duration of PBH domination.
Besides these low-frequency peculiarities, a PBH+superstrings (metastable strings) system provides a characteristic high-frequency GW spectrum constituting gravitons as shown in the bottom panels of Fig.7 and 8, where the GW strain $h_{c}$ is related to $\Omega_{\rm GW}$ as
@@
Because Eq.(IV.5) characterizes the GW background, for the allowed mass ranges of PBH presented in Sec.III, a considerable GW background constituting gravitons can be obtained within a constrained frequency range spanning approximately, only a decade. Although the proposed high-frequency GW detectors [143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153] are not presently competitive with the $\Delta N_{\rm eff}$ bound from Planck [154] (according to Ref.[155], the electromagnetic resonant cavities can potentially reach below the dark radiation projections around $f\sim 10^{8}$ Hz; the black oscillating curve in Fig.7 and Fig.8), optimistically, in the future, graviton background may be detected. In any case, these marked spectral features in the total GW background are described with three characteristic frequencies $f_{\rm brk}$, $f_{\rm peak,I}$, and $f_{\rm peak,g}$, making the PBH + string scenario more appealing than an ordinary EMD + string case.
@@ IV.2 Implication on a realistic BSM model of leptogenesis and super-heavy dark matter
where we have assumed leptogenesis is dominated by the decay of $N_{1}$ so that $M_{1}\sim T_{\rm lepto}$. Eq.(IV.8) can be shown to be accurate, matching the numerical results obtained from Eq.(III.1),(III.2) and the modified leptogenesis equations in the presence of PBHs [52]. The PBHs can also populate the decaying RHNs to produce lepton asymmetry, but if Eq.(IV.7) is to be satisfied, such PBHs evaporate after Sphaleron freeze-out temperature, and the lepton asymmetry does not get processed to BAU. Therefore we discard this option.
@@
$\mu=\pi M_{\rm DM}^{2}~{}f_{\rm DM}^{-2}h\left(\lambda,g^{\prime}\right),$ (IV.9)
@@
With this relation, we now redo the PTA fit for metastable strings performed in the previous section. The result is shown in Fig.10. Note that because of this extra condition in Eq.(IV.10), this $U(1)_{B-L}^{\rm PBH}$ model fits the PTA data for an extremely constrained range of PBH masses and produces GW signals belonging only to Category A. Therefore any non-observation of SGWB in the LVK-D would rule out the model/the best parameter choice of the model. Besides, because the PBH mass window is narrow, the other GW signals, e.g., from density fluctuations plus in the form of gravitons are peaked in a stringent range in frequencies. In an upcoming publication, we shall provide a detailed discussion of the model for more general parameter space with all possible GW signatures.
@@ V Conclusion
An abundant presence of ultralight PBHs leading to their domination on energy densities in the early universe could produce plenty of observable signatures specifically in the form of gravitational radiation. In addition to producing direct GW background, e.g., in the form of gravitons, a PBH domination gets imprinted on the spectral shape of GWs originating from other independent sources. Such a spectral shape reconstruction with the GW detectors and studying complementary GW signals originating directly from the PBHs could be an interesting way to probe the properties of ultralight PBHs. Here we study the impact of PBH domination on the GWs emitted from cosmic superstrings and metastable strings which are speculated to be among the sources of nHz GWs observed by the PTAs. Such an interpretation might become perilous, as a GW background originating from cosmic strings spans a wide range in frequencies thereby contradicting the non-observation of stochastic GW background by LIGO dealing with comparable signal strength. In fact, the LIGO constraint rules out a significant string-parameter space favoured by the recent PTA data. A PBH-dominated early universe, however, affects the propagation of such GWs to reduce the amplitude at the high frequencies thereby evading the LIGO constraints. Assuming monochromatic and non-rotating PBHs we work out the mass range of such ultralight PBHs which makes the string-radiated GWs consistent with LIGO bound and the PTA data. We show that PBHs with mass in the range $M_{\rm BH}\in\rm\left[10^{6}g,10^{8}g\right]$ would produce such viable signals that can be tested in the mid-band detectors such as LISA, high-frequency detectors such as LVK-D plus ET with a high signal-to-noise ratio. We discuss why a PBH-domination is distinguishable from any other early matter-domination scenarios that reduce the amplitude of string-radiated GWs at higher frequencies. In this context, we discuss a BSM model based on $U(1)_{B-L}$-seesaw providing neutrino mass, leptogenesis, super-heavy dark matter, and metastable cosmic strings. We show how a PBH-cosmic string scenario could be extremely predictive in probing unexplored parameter space of such BSM scenarios.
@@ Appendix A Superstrings and Metastable strings fit to the NANOGrav data
Figure 11: Triangular plot on the left (right) from the MCMC runs for superstrings (metastable strings) with the NANOGrav 15-year dataset using PTArcade[164] in the ceffyl mode[165].